One side of a square is 3x + 2 feet long. Another side is 5x - 8 feet long. How long is each side of the square?
Answer by ewatrrr(24785)

 
Hi,
One side of a square is 3x + 2 feet long. Another side is 5x - 8 feet long.
How long is each side of the square? Note: all sides of a square are of equal length
3x + 2 = 5x - 8
Solving for x
10 = 2x
5 = x
And 3x+2 = 17, the length of each side of the square
CHECKING our Answer***
(Note: 5x-8 = 17 as well)